Hi will you please explain to me how to determine whether each relation is linear. Here is one problem:
y=-4

Answer by fractalier(6550)

The word "linear" means a straight line. Normally any relation or function that is a first power in x is linear. Examples are y = 3x + 8, f(x) = 3/4 x + 2, y = x - 7, etc.
This case, y = -4, is a special case called a constant function...a horizontal line, in fact, without x at all.
It can be written as y = 0x - 4 to show you that it is also a linear relation.